OpenStudy (anonymous):

If f(x) = 2x 2 + 3, find f(3).

OpenStudy (solomonzelman):

\(\large\color{black}{ \bf f(x) = 2x^2 + 3 }\) \(\large\color{black}{ \bf f(-1) = 2(-1)^2 + 3~~~~~=2+3~~~~~=5 }\) \(\large\color{black}{ \bf f(4) = 2(4)^2 + 3~~~~~=32+3~~~~~=37 }\) \(\large\color{black}{ \bf f(3) = 2(3)^2 + 3=~~? }\)